### v1

# Intelligent Design 4.0: Paradigm Evolution Toward the Agentic AI Era

Abstract
Research and practice in Intelligent Design (ID) have significantly enhanced engineering innovation, efficiency, quality, and productivity over recent decades, fundamentally reshaping how engineering designers think, behave, and interact with design processes. The recent emergence of Foundation Models (FMs), particularly Large Language Models (LLMs), has demonstrated general knowledge-based reasoning capabilities, and open new paths and avenues for further transformation in engineering design. In this context, this paper introduces Intelligent Design 4.0 (ID 4.0) as an emerging paradigm empowered by agentic AI systems. We review the historical evolution of ID across four distinct stages: rule-based expert systems, task-specific machine learning models, large-scale foundation AI models, and the recent emerging paradigm of multi-agent collaboration. We propose a conceptual framework for ID 4.0 and discuss its potential to support end-to-end automation of engineering design processes through coordinated, autonomous multi-agent-based systems. Furthermore, we discuss future perspectives to enhance and fully realize ID 4.0’s potential, including more complex design scenarios, more practical design implementations, novel agent coordination mechanisms, and autonomous design goal-setting with better human value alignment. In sum, these insights lay a foundation for advancing Intelligent Design toward greater adaptivity, autonomy, and effectiveness in addressing increasingly complex design challenges.
Keywords Intelligent Design  $\cdot$ Engineering Design  $\cdot$ Design Automation  $\cdot$ Artificial Intelligence  $\cdot$ Large Language Model  $\cdot$ Multi-Agent System
1 Introduction
Design is a central activity in engineering [1, 2]. It serves as the engine for transforming multi-level needs and constraints into functional, manufacturable, and impactful products or solutions [3]. Historically grounded in expert heuristics and iterative human judgment, design has evolved with the development of computational tools, including drafting and modeling software such as Computer-Aided Design (CAD), simulation and optimization toolkits in Computer-Aided Engineering (CAE), and ideation support tools in Computer-Aided Ideation (CAI). As design problems become more complex, interdisciplinary, and dynamic, traditional design paradigms face limitations in terms of scalability and adaptivity.
Intelligent Design (ID) has emerged as a cross-disciplinary field at the intersection of engineering and artificial intelligence, with the goal of embedding reasoning, learning, and automation into design processes. ID includes a broad set of methods, tools, and systems that augment or automate design tasks through artificial intelligence. ID shares conceptual and methodological overlap with terms such as Design Automation [4] and Data-Driven Design [5], while covering a more inclusive scope that unifies all forms of computational intelligence aimed at supporting the engineering design process. Over successive technological waves, from rule-based or knowledge-based engineering systems [6], data-driven inference [7], to deep generative models [8], ID has reshaped how design is represented, explored, and realized.
More recently, the emergence of large-scale foundation models (e.g., Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs)) pretrained on vast corpora of multimodal data and refined through supervised fine-tuning and reinforcement learning, has demonstrated broad general-purpose capabilities [9, 10, 11]. These models have shown promise across diverse domains, including scientific discovery and engineering practice [12, 13]. In engineering design, recent research has explored their application in design tasks such as requirement extraction [14], design ideation [15], design modeling [16], and design optimization [17]. However, these efforts largely support only one or several isolated stages of the overall design process and continue to rely heavily on human efforts for orchestration and decision-making. As these LLM-based and MLLM-based design methods and tools increasingly exhibit competence, they pave the way for a new paradigm of more autonomous and agentic AI systems.
Herein, this paper aims to position such a paradigm, which we term Intelligent Design 4.0 (ID 4.0). We retrospectively analyze the evolution of Intelligent Design across four developmental stages, present a conceptual framework for multi-agent design collaboration, and discuss several future perspectives for advancing engineering design in the era of increasingly general and collaborative AI agent systems.
Figure 1: The Four-Stage Evolution of the Intelligent Design Paradigm
2 The Evolution of Intelligent Design Paradigms
Over the past decades, Intelligent Design has evolved from static, manually encoded expert systems to increasingly adaptive, data-driven, and now potentially autonomous systems powered by multiple AI agents. This paradigm evolution is mainly driven by three essential shifts:
  1. (1)
    The expansion of the design-related knowledge of AI, from limited and domain-specific datasets to vast and cross-domain data sources.
  2. (2)
    The progression of AI capabilities from task-specific assistance to general-purpose reasoning and decision-making.
  3. (3)
    The transition from human-in-the-loop control toward partially or fully autonomous design agents.
These shifts are redefining how design problems are framed, how solution spaces are explored, and how final design decisions are made. We identify four major stages in the trajectory of ID’s evolution, as illustrated in Figure 1.
3 Intelligent Design 1.0: Rule-Based Expert Systems
The first generation of Intelligent Design systems emerged in the 1980s-1990s, grounded in symbolic AI and expert systems. These rule-based systems aimed to capture and formalize expert knowledge into machine-readable formats for reuse. The underlying assumption was that design knowledge could be articulated explicitly and embedded into deterministic and heuristic rules [18].
A foundational aspect of ID 1.0 was the development of design ontologies and representation frameworks, such as the Function-Behavior-Structure (FBS) [19], SAPPhIRE [20], Functional basis [21], and Environment-Based Design (EBD) [22] ontologies. These frameworks enabled structured reasoning about design intent, function realization, and system behavior, facilitating design knowledge reuse [23]. In terms of methodology, ID 1.0 systems were often built upon structured design methods such as TRIZ, C-K Theory, Quality Function Deployment (QFD), and Axiomatic Design, which provided prescriptive strategies for exploring and evaluating solution spaces [24, 25, 26, 27, 28]. In parallel, early CAD software systems (e.g., AutoCAD, SolidWorks, Pro/E) also played a foundational role in Intelligent Design by automating repetitive operations and modeling tasks. While not traditionally categorized as AI, these tools significantly reduced manual effort and laid the groundwork for embedding structured logic into design workflows.
Figure 2: Representative ID 1.0 Systems: (A) DANE [29], a design by analogy to nature engine; (B) INNOGPS [30], a computer-aided system for design ideation and exploration
The computational realization of ID 1.0 was exemplified in early Knowledge-Based Engineering (KBE) and Case-Based Reasoning (CBR) systems, such as KRITIK [31], Argo [32], DANE [29], and InnoGPS [30], which supported analogy-based design through knowledge retrieval and re-organization. In addition, Design Structure Matrix (DSM) methods and relevant tools were also widely used to represent and manage interdependencies in complex engineering systems [33]. While effective in specific domains, these systems were inherently static, limited by their reliance on manually curated knowledge and rule sets. Figure 2 presents representative ID 1.0 systems that demonstrate early efforts in rule-based and knowledge-based design support.
4 Intelligent Design 2.0: Machine Learning-Driven Design
With the rise of computational power and the advancement of statistical learning techniques, especially the deep neural networks, ID 2.0 emerged with the widespread adoption of machine learning (ML) and deep learning (DL) methods [34]. ID 2.0 systems can learn correlations and latent structures from small-to-medium-scale datasets curated to specific design tasks, presenting a shift from expert-derived rules to data-driven computational reasoning.
One major direction of ID 2.0 was the development of systems for design information retrieval and design inspiration recommendation. Initial work leveraged structured textual data, such as thesaurus-based retrieval of informal design descriptions [35]. Later research extended these capabilities to include technical documents [36], images [37], sketches [38], and multimodal datasets that integrated textual, visual, and 3D models to enhance the discovery of analogous or novel design stimuli [39].
For design method recommendation, ML and DL algorithms were employed to identify relevant design principles or process heuristics. For example, supervised models have been trained to recommend suitable methods based on problem representations and user intent [40]. Another study mined latent patterns in designers’ workflows using Hidden Markov Models (HMM), which captured the sequential structure of human decision-making in design tasks for new design generation [41].
In recent years, Generative AI (GenAI) models introduced a new form of machine creativity [8]. Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), and their derivatives were developed and adopted to generate novel design candidates under functional, structural, and performance constraints [8]. These models have been applied on various design cases, such as wheels [42], bicycles [43], airfoils [44], and vehicles [45], demonstrating the power of GenAI methods in constrained design synthesis.
Additionally, techniques such as Natural Language Processing (NLP) [46], Semantic Networks (SN) [47], and Multi-Modal Machine Learning [48] also expanded the ability to model and reason over complex design-related tasks. By leveraging various forms of structured and unstructured data [7, 49], ID 2.0 systems enabled more intelligent representation, search, recommendation, and generation, ultimately improving design efficiency, creativity, and decision support across a wide range of tasks. Figure 3 presents representative ID 2.0 systems, which leverage ML/DL techniques to support specific tasks in the engineering design workflow.
Figure 3: Representative ID 2.0 Systems: (A) DL-based design stimuli search system that retrieves analogous visual designs [39]; (B) DL-based linkage design generation system that synthesizes mechanical linkages [50]
5 Intelligent Design 3.0: Foundation Model-Empowered Design
The third stage in the evolution of Intelligent Design is driven by the introduction of large-scale foundation AI models into engineering design workflows, including LLMs and MLLMs. These models are pretrained on vast multimodal corpora and exhibit emergent abilities such as in-context learning, instruction following, and step-by-step reasoning [51, 52]. In engineering design, recent studies have systematically evaluated their applicability and limitations on design-related tasks, demonstrating significant potential to support various stages of the design process [53, 54, 55, 56, 57].
Despite their remarkable reasoning power and adaptability, foundation AI models face critical challenges, such as hallucination, reasoning inconsistency, and the lack of verifiability [58]. To mitigate these limitations, researchers have proposed new methodologies aimed at better aligning model behavior with design-specific requirements. These include:
  1. (1)
    defining explicit task workflows to guide the reasoning steps of LLMs in requirement analysis, design generation, and evaluation;
  2. (2)
    combining lightweight task-specific models with general-purpose LLMs to improve output fidelity and domain relevance;
  3. (3)
    integrating retrieval-augmented generation (RAG) or domain-specific fine-tuning to reinforce factual grounding and context consistency.
With these strategies, researchers have developed LLM-based and MLLM-based approaches to enhance one or multiple stages of the design processes, including requirement analysis [14, 59, 60], concept generation [15, 56, 61, 62], system construction [63], CAD modeling and editing [16, 64, 65], and design optimization [17, 66]. These efforts reflect a shift from using foundation AI models as general copilots toward positioning them as adaptable, architecture-integrated components within intelligent design systems. Compared to ID 2.0, where models largely automated decisions or learned patterns from labeled data, ID 3.0 systems are increasingly capable of complex reasoning and reflective decision-making. Figure 4 presents representative ID 3.0 systems that leverage large-scale foundation models for conceptual ideation and detailed CAD modeling.
Figure 4: Representative ID 3.0 Systems: (A) ZOO [65], an LLM-based Text-to-CAD system; (B) AutoTRIZ [15], an artificial ideation system with TRIZ and LLMs
6 Intelligent Design 4.0: Agentic and Multi-Agent AI Systems
While LLM-driven and MLLM-driven methods presents advanced capabilities to design, a single foundation model remains insufficient to handle the complexity of the full stack of engineering design. The whole design process spans a wide range of cognitive, computational, domain-specific subtasks. The design methods and tools of ID 3.0 still require human oversight for task decomposition, context switching, and coordination [67]. In addition, recent studies have also revealed intrinsic capability variations across different models [68]. For instance, some LLMs such as Claude models excel at code generation, while others like GPT models demonstrate stronger performance in text generation [9, 10]. These observations suggest that the complex design tasks, are better served not by a single foundation model, but by an ensemble of specialized, collaborative AI agents, each fine-tuned or prompted for particular functions [69]. Figure 5 illustrates a fundamental framework of an AI agent. Powered by LLMs or MLLMs, AI agents can be defined as autonomous systems that can perceive its environment, break down complex tasks, evoke tool uses, learn from interactions (with both human users or other agents), and take actions toward achieving defined or evolving goals [69].
Figure 5: The fundamental framework of an AI Agent
Building upon the foundation of AI agents, the paradigm of ID 4.0 envisions multi-agent system composed of autonomous, design task-specialized agents that coordinate and collaborate to complete complex, end-to-end design processes. We propose a conceptual framework for implementing ID 4.0, as illustrated in Figure 6.
Figure 6: The Multi-Agent-based ID 4.0 Paradigm
In Figure 6, the left panel outlines a five-stage engineering design workflow adapted from [3]: Task Clarification, Concept Design, Embodiment Design, Detailed Design, and Design Optimization. These stages represent a typical iterative design loop, where solutions are progressively refined and evaluated. The right panel maps this framework onto a multi-agent AI system, where each design stage is managed by a corresponding stage-level agent:
  • Requirement Analysis Agent interprets user inputs, extracts design goals and constraints, and decomposes high-level tasks into sub-problems for downstream agents.
  • Concept Generation Agent explores early-stage solutions using ideation tools and creative reasoning, often invoking supporting agents like Brainstorming and Empathy agents for divergent thinking.
  • Embodiment Generation Agent translates selected concepts into more concrete forms, coordinating with modeling and simulation agents to assess feasibility and functionality.
  • Detail Modeling Agent refines the geometry and specifications of the design, producing manufacturable and parametric models in collaboration with supporting agents like CAD agents and design databases.
  • Design Optimization Agent applies generative and simulation-based optimization to fine-tune performance against multiple objectives or constraints.
These stage-level design agents are not static modules, but interactive and adaptive components. Each stage-level agent may operate as an individual LLM/MLLM-powered agent or as a coordinated group of agents sharing the same objective. Throughout the design process, each agent can dynamically take the lead in decomposing tasks, triggering downstream agents, and coordinating others as needed. To support complex operations, each stage-level agent accesses a shared pool of functional agents that provide domain-general capabilities across the design workflow:
  • Web Search Agent retrieves up-to-date external knowledge and references from the internet.
  • Database Agent accesses structured domain-specific datasets or internal databases to support data-informed decisions.
  • Empathy Agent infers user intent and emotional context to align design proposals with users’ expectations.
  • Brainstorming Agent facilitates divergent thinking to generate alternative concepts and creative directions.
  • CAD Modeling Agent synthesizes 2D and 3D geometries and integrates parametric constraints.
  • CAE Simulation Agent conducts multi-physics simulations (e.g., structural, thermal, and fluid) to evaluate functional performance of designs.
  • Requirement Check Agent verifies compliance with user-defined goals, regulatory constraints, and quality criteria.
These functional agents are not merely passive tools like external softwires or APIs. They are also equipped with capabilities for planning, adaptation, and system information sharing, enabling more intelligent and coordinated collaboration with stage-level agents. While Figure 6 illustrates a representative set of agents, real-world design environments may incorporate additional or domain-specific agents to meet specialized functional needs.
The entire multi-agent system interacts with human designers (initiate tasks and receive design outputs) through a centralized communication interface. Within the system, all agents can access external tools and services via standardized protocols [70]. Additionally, agents operate within a shared information environment, comprising both short-term and long-term memory, which support cross-agent coordination and cumulative learning throughout the design process.
This agentic framework enables iterative collaboration and distributed task execution across design stages. For example, the Concept Generation Agent may invoke the Brainstorming Agent and Empathy Agent to explore early-stage ideas, and then transfer selected candidates to the CAD Modeling Agent and CAE Simulation Agent for embodiment and performance evaluation. The Requirement Check Agent ensures that all design outputs remain aligned with user-defined goals and a range of desired system properties such as quality, reliability, and maintainability that often emerge during operational use [71]. In recent, several studies have explored the application of AI agents in design practice, validating on engineering problems such as aerodynamic vehicle design and mechatronic system design [72, 73, 74, 75]. By reducing the dependence on human intervention for coordination and design decision-making, ID 4.0 has the potential to fundamentally transform engineering design creativity and productivity, enabling continuous and autonomous design processes without being limited by human time, effort, or expertise.
7 Challenges and Opportunities Ahead for ID 4.0
As ID 4.0 evolves, its potential lies not only in automating existing design workflows, but in reimagining the nature of intelligence, autonomy, and collaboration in engineering. It brings a shift from static processes to dynamic ecosystems of learning, adaptation, and coordination. While recent studies have validated early prototypes of multi-agent design systems powered by foundation AI models [72, 73, 74, 75], some important open questions remain, particularly regarding how such systems can be scaled, grounded, coordinated, and goal-setting in increasingly complex and practical design contexts.
(1) Toward More Complex Design: Multi-scale and Multi-physics Co-design
As engineering design problems increase in complexity, they increasingly require integrated reasoning across disparate physical scales (e.g., nano to macro) and physics domains (e.g., structural, thermal, fluid, electromagnetic) [76]. Complex products such as aerospace systems, robotics, and biomedical devices often involve tight coupling between microstructural properties, component-level geometries, control strategies, and performance in dynamic environments [77, 78, 79].
Traditional design methods often treat these physical domains independently, requiring manual iteration across disparate software tools and collaboration among multiple disciplinary teams. Within the ID 4.0 paradigm, specialized agents can be orchestrated to operate at different modeling scales, enabling coherent and system-level co-design. For example, a material-level simulation agent could inform a structural optimization agent via shared intermediate representations, facilitating more comprehensive performance integration.
Despite this potential, current foundation models still lack accurate embedded knowledge in fields such as nonlinear mechanics, heat transfer, or optics [80]. Their reasoning is frequently statistical, not governed by physical laws. Closing this gap may require hybrid systems that integrate MLLM-powered agents with symbolic physics engines or world models [81]. The advancement of such hybrid systems could yield cross-scale design agents capable of reasoning from atomic structures to macro-level artifacts within multi-physics design environments.
(2) Toward More Practical Design: Context-aware Design with Real-world Constraints
While ID 4.0 have made significant achievements in generating functional and innovative design solutions, these outcomes must ultimately align with the practical realities of implementation. In real-world engineering, design decisions must account for a variety of non-ideal constraints, including manufacturing capabilities, supply chain limitations, material availability, cost structures, and environmental regulations [82, 83, 84]. A design that performs optimally in simulation may still fail if it cannot be manufactured efficiently or sourced reliably. Moreover, design outputs, especially when conceived as complex systems, must consider a set of quality attributes (commonly referred to as “ilities” [85]), such as reliability, maintainability, scalability, and quality. These attributes are critical to ensure that the system can deliver sustained value to stakeholders and users over its lifecycle [86].
To move toward more practically deployable design intelligence, ID 4.0 must become context-aware, embedding real-world data into the reasoning loop. This requires integrating agents with external systems or data sources such as manufacturing databases, supplier networks, market analytics platforms, and internal databases in organizations like Enterprise Resource Planning (ERP) systems. Through such integration, agents can dynamically retrieve and process contextual information, such as material availability, production tolerances, cost trends, carbon footprints, and supply chain risks.
When multiple constraints conflict, design agents must balance trade-offs through principled methods. In this case, agents may involve multi-objective optimization to reason over competing criteria, such as cost vs. performance or speed vs. sustainability. These optimization components can function as internal capabilities or external tools invoked via standardized protocols.
(3) Toward More Effective Agent Collaboration
Current research in multi-agent intelligent design systems often relies on a rigidly defined design process to orchestrate collaboration among agents [72]. While such structured workflows are valuable for feasibility studies and benchmarking, they may be limited in more open-ended design scenarios, where creative exploration and emergent strategy are essential.
To move beyond these limitations, ID 4.0 demands more flexible and dynamic coordination mechanisms, potentially inspired by concepts from self-organizing systems or multi-agent negotiation [87]. This remains an open question for design practice: how can design agent teams autonomously allocate tasks, exchange knowledge, and evolve their collaboration strategies effectively?
Furthermore, as foundation models continue to grow in capacity, the boundary between the multi-agent framework and the single “super-agents” may become blurred [88]. This raises critical questions: At what point could a single model effectively manage the full stack of an end-to-end design workflow? And how should we balance the generality and power of such models against the modularity, interpretability, and fault isolation offered by multi-agent systems? These questions lie at the heart of the orchestration problem in ID 4.0: what is the optimal architecture for intelligent collaboration - distributed agents with explicit communication and specialization, or unified agents with general-purpose capabilities? Future research and practice must explore hybrid architectural strategies that can dynamically evolve with advancements in model capabilities, balancing the strengths of specialization and generality within intelligent design systems.
(4) Toward New Forms of Goal-setting
Engineering design has long been a human-centered activity, with goals explicitly specified by users based on functional requirements, stakeholder needs, and contextual constraints [89]. Even under the ID 4.0 paradigm, human designers typically serve as the ultimate arbiters of what constitutes a desirable outcome [60]. However, as agentic AI systems grow more capable of reasoning, interacting, and learning from context, a new possibility emerges: Can AI agents begin to formulate goals themselves? For example, in response to vague or under-specified prompts, an ID 4.0 system might extrapolate user intent, identify broader system-level opportunities, or suggest higher-level objectives that were previously unconsidered.
The shift from human-defined goals to machine-generated goals may present a fundamental change in how design is conceived and directed [90]. It challenges the long-standing assumption that human intention is the sole source of meaningful goals. If agents are empowered to redefine the design problem or elevate the abstraction level of design targets, they may unlock higher-order creativity and autonomous problem discovery. In doing so, ID 4.0 could evolve from a tool for goal execution into a partner in ideation, capable of augmenting or even co-driving the innovation process.
At the same time, the shift toward machine-defined goals raises fundamental ethical and legal questions [91]. Who is accountable for a machine-generated goal? How do we ensure that such goals respect cultural, environmental, or organizational norms? And what does authorship mean when systems contribute not only as problem-solvers but also as problem-setters?
To better align ID 4.0 systems with human values, we need alignment on two fronts: technically, through transparent reasoning processes and human-in-the-loop validation; and socially, by embedding ethical norms and stakeholder priorities into system design. In addition, relevant policies should ensure that system actions are traceable, responsibilities are well-defined, and human intervention remains possible when needed.
8 Conclusion
Intelligent Design 4.0 represents a critical inflection point in the ongoing evolution of design intelligence, shifting from data-driven assistance to multi-agent autonomy empowered by foundation AI models. In this paper, we reviewed the paradigm evolution of Intelligent Design through four developmental stages, and presented a multi-agent-based framework for ID 4.0. Beyond reviewing past developments, we have outlined future directions and open challenges toward a more scalable, practical, effective design paradigm. We hope this work offers both conceptual foundations and actionable perspectives to guide researchers, designers, and developers in advancing the next generation of intelligent design systems.
Disclosure statement
No potential conflict of interest was reported by the author(s).